OpenStudy (fools101):

can anyone help with Inverse Functions and who to brake them down have problems with is!!! f(x) = 2x2 - 1

OpenStudy (across):

Let\[f(x)\]be a bijection such that\[f:\mathbb{R}\to\mathbb{R}.\]The function can then have an inverse, denoted as\[f^{-1}(x),\]such that\[f^{-1}(f(x))=x.\]E.g., to find the inverse of\[f(x)=y=2x^2-1,\]you swap the variables and solve for y:\[x=2y^2-1.\]Can you solve for y?

OpenStudy (fools101):

yes we can use y?

OpenStudy (fools101):

can we?

OpenStudy (fools101):

wow that was confuse I'm still confuse :(

OpenStudy (across):

\[x=2y^2-1.\]How can you solve for y? You add 1 to both sides:\[x+1=2y^2,\]then you divide both sides by 2:\[\frac{x+1}{2}=y^2,\]and finally, you take the square root of both sides:\[y=\pm\sqrt{\frac{x+1}{2}}.\]However, note that this is not a function.

OpenStudy (fools101):

f(x) = 2x2 - 1 so i change it to a Y?

OpenStudy (fools101):

so it would be F(Y) = 2x2-1?

OpenStudy (across):

If you have a function\[y=f(x),\]you swap the variables and solve for y:\[x=f(y).\]

OpenStudy (fools101):

oh but i don't for the Answer it show this F^-1 for all the Answer but my problem is f(x) = 2x2 - 1 idk who to brake it down

OpenStudy (across):

I already did it for you up there.

OpenStudy (fools101):

like this for Example: f(x) = (x - 1)2 - 5 what i did for this one was like Y = (x-1)^2 -5 Y = (x-1)^2 - 5 X = (y - 1)^2 -5 mathml equation = mathml equation mathml equation

OpenStudy (fools101):

\[\sqrt{X +5} = \sqrt{Y- 1} \sqrt{X+5}\] and X+ 5 is my Answer.

OpenStudy (fools101):

while i know but non of my Answers have nothing to do with y?

OpenStudy (across):

Okay, let me work that one out step by step: We have that\[y=(x-1)^{2}-5.\]To find the inverse of this function, we first swap variables:\[x=(y-1)^{2}-5.\]Now that we have swapped variables, we want to solve for y alone. First, we add 5 to both sides:\[x+5=(y-1)^{2}.\]Next, we take the square root of both sides:\[\pm\sqrt{x+5}=y-1.\]Finally, we add 1 to both sides:\[\pm\sqrt{x+5}+1=y.\]Therefore, we see that\[y=\pm\sqrt{x+5}+1,\]and we have found the function's inverse.

OpenStudy (fools101):

Awww know that one but who would this work f(x) = 2x^2 - 1

OpenStudy (fools101):

how*

OpenStudy (across):

Again, I did it for you up there...

OpenStudy (fools101):

ohh kk that make no since but thank u Across for trying to help!